# Benedict Quartey, Ifrah Idrees, Jason Xinyu Liu, Stefanie Tellex

### v2

@@ I INTRODUCTION
We used $\lambda$ to benchmark several models and establish baseline performance for long-horizon MoMa tasks. Behavior cloning (BC) models trained on the dataset showed low success rates, $2.7\%$ and $5.1\%$, indicating data inefficiency while confirming the dataset’s ability to support learning. A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an $44.4\%$ success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems in the future. These results demonstrate $\lambda$’s effectiveness in highlighting data efficiency limitations of current models and guiding the future development of improved architectures for indoor robotics.
@@ IV-A Baseline Models
To evaluate the $\lambda$ and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic models. Given the growing popularity of imitation learning (IL), particularly BC, we benchmark a state-of-the-art (SOTA) BC model. These models provide insight into how data-efficient current BC algorithms are by being tested on $\lambda$, validating that the dataset supports meaningful learning by demonstrating that models can achieve measurable yet limited success. Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP). This system, evaluated zero-shot without requiring training, investigates whether non-IL methods can achieve greater data efficiency while achieving comparable or superior performance, offering guidance for future development of data-efficient MoMa methodologies.
@@
TABLE II: Data Splits. Splitting of the simulation dataset of 521 demonstrations into training, validation, and held-out test subsets for the Scene Generalization and Task Generalization experiments. For Scene Gen., a stratified sampling approach was employed to account for variations in the number of demonstrations per scene (details in Section IV-B). The “Scenes” row indicates the number of scenes contributing to each subset of demonstrations.
@@ V-A Results
Task Generalization. The success rates for the Task Generalization experiment for RT-1, as presented in Table IV, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment. This improvement is likely attributed to the model having access to all the scenes during training, enabling it to better generalize to new tasks within already familiar environments.
@@
LIMP. The results presented thus far indicate that learning-based models, such as RT-1, face significant challenges in addressing the complexities of long-horizon MoMa tasks. To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
@@ V-B Ablations
TABLE III: Scene Generalization results. Success rates for the Scene Generalization experiment on models trained from scratch. The success rate is computed by averaging the scores of four subtasks per task, with each successful subtask contributing one point, and then averaging across all tasks. The subtasks are: navigation to the target object, grasping the target object, navigation to the goal with the target object, and placing the target object at the goal. $\pm$ indicates the standard deviation across all evaluated tasks. “Avg” indicates the average of all the scenes.
@@
TABLE IV: Task Generalization results. Success rates for Task Generalization experiments, encompassing both training from scratch and fine-tuning of pretrained parameters. The computation is the same as Table III. “Seen” refers to evaluation on the training subset, while “Unseen” refers to evaluation on the held-out test subset. Denotes zero-shot evaluation with no training on a mix of simulation and real-world tasks.
Impact of Architecture. To determine if data inefficiency stems from the architecture used, we conduct a model comparison experiment by evaluating a different architecture. We developed a custom LSTM-based model by replacing RT-1’s Transformer with an LSTM while retaining the same encoders for a consistent and isolated comparison, as LSTMs have been previously used for sequential robotic tasks. The LSTM variant is evaluated under the task generalization setup using subset experiments with $25\%$, $50\%$, $75\%$, and $100\%$ of the dataset, enabling us to analyze how its performance scales with increasing data. By comparing the full dataset results and the performance rate of change (RoC) against RT-1, we aim to determine whether data inefficiency is tied to architectural limitations or if the issue lies elsewhere.
@@ VI Conclusion
To evaluate the benchmark, we tested multiple models, including RT-1 and LIMP, to assess their performance and data efficiency on these complex tasks. Learning-based models such as RT-1 demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks. In contrast, the neuro-symbolic system LIMP significantly outperformed RT-1, showcasing its potential as a promising direction for the development of future mobile manipulation systems capable of addressing the challenges posed by $\lambda$’s tasks. We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.
@@
ACKNOWLEDGMENT
This work is supported by ONR under grant award numbers N00014-22-1-2592 and N00014-23-1-2794, NSF under grant award number CNS-2150184, and with support from Amazon Robotics. We also thank Aryan Singh, George Chemmala, Ziyi Yang, David Paulius, Ivy He, Lakshita Dodeja, Mingxi Jia, Benned Hedegaard, Thao Nguyen, Selena Williams, Tuluhan Akbulut, Chaerin Min, and George Konidaris for their help in various phases of work.

### v3

@@ I INTRODUCTION
We used $\lambda$ to benchmark several models and establish baseline performance for long-horizon MoMa tasks. Behavior cloning (BC) models trained on the dataset showed low success rates, such as $2.7\%$ and $5.1\%$, indicating data inefficiency while confirming the dataset’s ability to support learning. A neuro-symbolic modular approach, combining foundation models with task and motion planning, achieved an $44.4\%$ success rate, significantly outperforming the BC models, illustrating the potential of such approaches to enable more data-efficient MoMa systems in the future. These results demonstrate $\lambda$’s effectiveness in highlighting data efficiency limitations of current models and guiding the future development of improved architectures for indoor robotics.
@@ IV-A Baseline Models
To evaluate the $\lambda$ and establish baselines for future benchmarking, we test both learning-based and neuro-symbolic models. Given the growing popularity of imitation learning (IL), particularly BC, we benchmark two state-of-the-art (SOTA) BC models. These models provide insight into how data-efficient current BC algorithms are by being tested on $\lambda$, validating that the dataset supports meaningful learning by demonstrating that models can achieve measurable yet limited success. Additionally, we benchmark a neuro-symbolic system that integrates foundation models with task and motion planning (TAMP). This system, evaluated zero-shot without requiring training, investigates whether non-IL methods can achieve greater data efficiency while achieving comparable or superior performance, offering guidance for future development of data-efficient MoMa methodologies.
@@
MotionGlot. MotionGlot (MG) [25] is designed to address the problem of varying action and observation dimensions in multi-embodiment models [5, 26]. Leveraging training principles from language models (LLMs), MG proposes a training procedure for motion generation across multiple embodiments as a next-token prediction problem. Here we adapt MG for the MoMa tasks, given a set of demonstration trajectories with image observation ($o_{i}$) action ($a_{i}$) pairs in the form of $\tau=\{(o_{0},a_{0}),(o_{1},a_{1}),...,(o_{T},a_{T}))\}$, MG learns a model to map a given observation to action. This process involves two stages: First, a VQ-VAE is employed to learn a discrete latent codebook for the images, forming the observation vocabulary, while discrete binning is applied to tokenize the actions. In the second stage, both the observation and action vocabularies are integrated with the text vocabulary from GPT-2 [27]. This unified vocabulary allows for action generation similar to text generation. Finally, the instruction-tuning template from MG is used for text command-based action generation.
@@
TABLE II: Data Splits. Splitting of the simulation dataset of 521 demonstrations into training, validation, and held-out test subsets for the Scene Generalization and Task Generalization experiments. The “Scenes” row indicates the number of scenes contributing to each subset of demonstrations.
@@ V-A Results
Task Generalization. The success rates for the Task Generalization experiment for RT-1 and MotionGlot, as presented in Table IV, are low overall, though they are slightly higher compared to the average performance in the Scene Generalization experiment. This improvement is likely attributed to the model having access to all the scenes during training, enabling it to better generalize to new tasks within already familiar environments. Additionally, while it is unsurprising that RT-1 outperformed MotionGlot, it is notable that the margin was relatively small, particularly given that MotionGlot was originally designed for motion generation rather than specifically for mobile manipulation.
@@
LIMP. The results presented thus far indicate that learning-based models, such as RT-1 and MotionGlot, face significant challenges in addressing the complexities of long-horizon MoMa tasks. To explore an alternative paradigm, we benchmarked the LIMP system on the Task Generalization experiment.
@@ V-B Ablations
TABLE III: Scene Generalization results. Success rates for the Scene Generalization experiment on models trained from scratch. The success rate is computed by averaging the scores of four subtasks per task, with each successful subtask contributing one point, and then averaging across all tasks. The subtasks are: navigation to the target object, grasping the target object, navigation to the goal with the target object, and placing the target object at the goal.
@@
TABLE IV: Task Generalization results. Success rates for Task Generalization experiments, encompassing both training from scratch and fine-tuning of pretrained parameters. “Seen” refers to evaluation on the training subset, while “Unseen” refers to evaluation on the held-out test subset. Denotes zero-shot evaluation with no training on a mix of simulation and real-world tasks.
Impact of Architecture. To determine if data inefficiency stems from the architecture used, we conduct a model comparison experiment by evaluating a different architecture. While MotionGlot was already benchmarked and provides some architectural comparison with RT-1, it is not sufficiently different since it also relies on a transformer architecture. To explore a more distinct alternative, we developed a custom LSTM-based model by replacing RT-1’s Transformer with an LSTM while retaining the same encoders for a consistent and isolated comparison, as LSTMs have been previously used for sequential robotic tasks. The LSTM variant is evaluated under the task generalization setup using subset experiments with $25\%$, $50\%$, $75\%$, and $100\%$ of the dataset, enabling us to analyze how its performance scales with increasing data. By comparing the full dataset results and the performance rate of change (RoC) against RT-1, we aim to determine whether data inefficiency is tied to architectural limitations or if the issue lies elsewhere.
@@ VI Conclusion
To evaluate the benchmark, we tested multiple models: RT-1, MotionGlot, and LIMP, to assess their performance and data efficiency on these complex tasks. Learning-based models (RT-1 and MotionGlot) demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to $\lambda$’s long-horizon tasks. In contrast, the neuro-symbolic system LIMP significantly outperformed RT-1 and MotionGlot, showcasing its potential as a promising direction for the development of future mobile manipulation systems capable of addressing the challenges posed by $\lambda$’s tasks. We hope that this benchmark will serve as a valuable resource for the robotics community, fostering the development of more advanced models and systems that can efficiently address daily household and workplace tasks.